OpenStudy (anonymous):

I need help x^3 + 5x^2 - x - 5

OpenStudy (anonymous):

I think the question is factoring this polynomial. So we have \[= x^2(x+5) - (x+5)\] \[= (x+5)(x^2-1)\] \[=(x+5)(x-1)(x+1)\]